# Computer-Assisted Proofs of Gap Solitons in Bose-Einstein Condensates

### v2

@@ 3 Solving the Boundary-Value Problem
$Y(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})$ =$\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\left\|A_{f}\begin{pmatrix}\bar{s}^{(2)}(-1)\\ L\bar{s}+\kappa\ T\operatorname{\mathit{f}}(\bar{s})\end{pmatrix}\right\|_{X_{\scriptscriptstyle{\mathsf{C}}}}+\lVert L^{-1}\prod\limits_{\scriptscriptstyle{(M,3M+1]}}^{\mathsf{\scriptscriptstyle{C}}}F(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\rVert_{X_{\scriptscriptstyle{\mathsf{C}}}}$ (62)
$+\lVert A_{f}\rVert_{B(X_{\scriptscriptstyle{\mathsf{C}}})}\max\left\{\right.|\bar{s}^{(1)}(1)-\bar{W}^{(1)}(\theta,\bar{\sigma})|+r_{\scriptscriptstyle{\mathsf{TF}}},|\bar{s}^{(2)}(1)-\bar{W}^{(2)}(\theta,\bar{\sigma})|+r_{\scriptscriptstyle{\mathsf{TF}}},|\bar{s}^{(3)}(-1)-1|,|\bar{s}^{(4)}(-1)|\left.\right\}.$
@@
$=\max\left\{\right.|\bar{s}^{(1)}(1)-W^{(1)}(\theta,\bar{\sigma})|,|\bar{s}^{(2)}(1)-W^{(2)}(\theta,\bar{\sigma})|,|\bar{s}^{(3)}(-1)-1|,|\bar{s}^{(4)}(-1)|\left.\right\}$
$\leq\max\left\{\right.|\bar{s}^{(1)}(1)-\bar{W}^{(1)}(\theta,\bar{\sigma})|+r_{\scriptscriptstyle{\mathsf{TF}}},|\bar{s}^{(2)}(1)-\bar{W}^{(2)}(\theta,\bar{\sigma})|+r_{\scriptscriptstyle{\mathsf{TF}}},|\bar{s}^{(3)}(-1)-1|,|\bar{s}^{(4)}(-1)|\left.\right\}.$
@@
$Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})$ =$\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\lVert\prod\limits^{\mathsf{\scriptscriptstyle{\mathbb{R}}}}\ +\prod\limits_{\scriptscriptstyle{[0,M]}}^{\mathsf{\scriptscriptstyle{C}}}-A_{f}DF(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\left(\prod\limits^{\mathsf{\scriptscriptstyle{\mathbb{R}}}}+\prod\limits_{\scriptscriptstyle{[0,3M+1]}}^{\mathsf{\scriptscriptstyle{C}}}\right)-A_{f}DB(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\prod\limits_{\scriptscriptstyle{[0,3M+1]}}^{\mathsf{\scriptscriptstyle{C}}}-A_{f}D\bar{B}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\prod\limits^{\mathsf{\scriptscriptstyle{\mathbb{R}}}}\rVert_{B(X_{\scriptscriptstyle{\mathsf{C}}})}$ (64)
@@
For the next term we apply an analogous result to Lemma 3. In this case, we consider a truncation up to mode $13M+$ 1 to account for the action of the operator $T$, which shifts the non-zero modes down by one. Indeed
@@
$\prod\limits_{\scriptscriptstyle{(3M+1,\infty)}}^{\mathsf{\scriptscriptstyle{C}}}-A_{\infty}DF(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\prod\limits_{\scriptscriptstyle{(3M+1,\infty)}}^{\mathsf{\scriptscriptstyle{C}}}&=-\kappa\ L^{-1}\prod\limits_{\scriptscriptstyle{(M,\infty)}}^{\mathsf{\scriptscriptstyle{C}}}TDf(\bar{s})\prod\limits_{\scriptscriptstyle{(3M+1,\infty)}}^{\mathsf{\scriptscriptstyle{C}}}\end{aligned}.$
@@
Moreover, we are supposing that $-1<\bar{\sigma}<1$ and $1\leq\nu$, hence
@@
$Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})$ =$\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,2\omega|\kappa|\left(\lVert A_{f}\rVert_{B(X_{\scriptscriptstyle{\mathsf{C}}})}+\frac{1}{2M}\right)\left(2|b|+6|c|\lVert\bar{s}^{(1)}_{\scriptscriptstyle{\mathsf{C}}}\rVert_{\scriptscriptstyle{\mathsf{C}}}+3|c|r\right)$ (69)
@@
For the term including the boundary conditions. We begin by considering an element $h=(\sigma_{h},s_{h})$ of $X_{\scriptscriptstyle{\mathsf{C}}}$ such that $\lVert h\lVert_{X_{\scriptscriptstyle{\mathsf{C}}}}\leq 1$ We have that
$\lVert\left(DB(x)-DB(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\right)h\rVert_{X_{\scriptscriptstyle{\mathsf{C}}}}$ =$=\lVert\prod\limits_{\scriptscriptstyle{\{0\}}}^{\mathsf{\scriptscriptstyle{C}}}\left(DB(x)-DB(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\right)h\rVert_{X_{\scriptscriptstyle{\mathsf{C}}}}$
@@
$Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})<1\quad\text{and}\quad Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{C}}}^{*})<min\left(\frac{(1-Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}}))^{2}}{2Y(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})},r_{\scriptscriptstyle{\mathsf{TF}}}^{*}\right)$
@@
$r_{\scriptscriptstyle{\mathsf{C}}}=\frac{1-Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})-\sqrt{(1-Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{C}}}^{*}))^{2}-2Y(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{C}}}^{*})}}{Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{C}}}^{*})}.$
@@ 4 Examples of Constructive Proofs of Existence of Gap Solitons
The Gross-Pitaevskii equation (2) with parameters $a=1.1025$, $b=0.55125$, and $c=-0.826875$ has a soliton solution $u:\mathbb{R}\to\mathbb{R}$, satisfying
@@
[…]
@@
[…]
$Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}})<1\quad\text{and}\quad Z_{2}(\bar{w},r_{\scriptscriptstyle{\mathsf{TF}}}^{*})<min\left(\frac{(1-Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}}))^{2}}{2Y(\bar{w},r_{\scriptscriptstyle{\mathsf{F}}})},r_{\scriptscriptstyle{\mathsf{TF}}}^{*}\right).$
@@
[…]
@@
$Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})<1\quad\text{and}\quad Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{C}}}^{*})<\min\!\Biggl{(}\frac{\bigl{(}1-Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})\bigr{)}^{2}}{2Y(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})},\;r_{\scriptscriptstyle{\mathsf{C}}}^{*}\Biggr{)}$
@@
The Gross-Pitaevskii equation (2) with parameters $a=1$, $b=-1$ and $c=1$, has a soliton solution.

### v3

@@ 3 Solving the Boundary-Value Problem
$Y(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})$ $\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\left\|A_{f}\begin{pmatrix}\bar{s}^{(2)}(-1)\\ L\bar{s}+\kappa\ T\operatorname{\mathit{f}}(\bar{s})\end{pmatrix}\right\|_{X_{\scriptscriptstyle{\mathsf{C}}}}+\lVert L^{-1}\prod\limits_{\scriptscriptstyle{(M,3M+1]}}^{\mathsf{\scriptscriptstyle{C}}}F(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\rVert_{X_{\scriptscriptstyle{\mathsf{C}}}}$ (66)
[…]
@@
[…]
@@
$Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})$ $\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\lVert\prod\limits^{\mathsf{\scriptscriptstyle{\mathbb{R}}}}\ +\prod\limits_{\scriptscriptstyle{[0,M]}}^{\mathsf{\scriptscriptstyle{C}}}-A_{f}DF(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\left(\prod\limits^{\mathsf{\scriptscriptstyle{\mathbb{R}}}}+\prod\limits_{\scriptscriptstyle{[0,3M+1]}}^{\mathsf{\scriptscriptstyle{C}}}\right)-A_{f}DB(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\prod\limits_{\scriptscriptstyle{[0,3M+1]}}^{\mathsf{\scriptscriptstyle{C}}}-A_{f}D\bar{B}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\prod\limits^{\mathsf{\scriptscriptstyle{\mathbb{R}}}}\rVert_{B(X_{\scriptscriptstyle{\mathsf{C}}})}$ (68)
@@
For the next term we apply an analogous result to Lemma 3. In this case, we consider a truncation up to mode $3M+1$ to account for the action of the operator $T$, which shifts the non-zero modes down by one. Indeed
@@
$\begin{aligned} \prod\limits_{\scriptscriptstyle{(3M+1,\infty)}}^{\mathsf{\scriptscriptstyle{C}}}-A_{\infty}DF(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\prod\limits_{\scriptscriptstyle{(3M+1,\infty)}}^{\mathsf{\scriptscriptstyle{C}}}&=-\kappa\ L^{-1}\prod\limits_{\scriptscriptstyle{(M,\infty)}}^{\mathsf{\scriptscriptstyle{C}}}TDf(\bar{s})\prod\limits_{\scriptscriptstyle{(3M+1,\infty)}}^{\mathsf{\scriptscriptstyle{C}}}\end{aligned}.$
@@
Moreover, we are supposing that $1<\bar{\sigma}<1$ and $1\leq\nu$, hence
@@
$Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})$ $\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,2\omega|\kappa|\left(\lVert A_{f}\rVert_{B(X_{\scriptscriptstyle{\mathsf{C}}})}+\frac{1}{2M}\right)\left(2|b|+6|c|\lVert\bar{s}^{(1)}_{\scriptscriptstyle{\mathsf{C}}}\rVert_{\scriptscriptstyle{\mathsf{C}}}+3|c|r\right)$ (73)
@@
For the term including the boundary conditions. We begin by considering an element $h=(\sigma_{h},s_{h})$ of $X_{\scriptscriptstyle{\mathsf{C}}}$ such that $\lVert h\lVert_{X_{\scriptscriptstyle{\mathsf{C}}}}\leq 1$ . We have that
$\lVert\left(DB(x)-DB(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\right)h\rVert_{X_{\scriptscriptstyle{\mathsf{C}}}}$ $=\lVert\prod\limits_{\scriptscriptstyle{\{0\}}}^{\mathsf{\scriptscriptstyle{C}}}\left(DB(x)-DB(\bar{x}_{\scriptscriptstyle{\mathsf{C}}})\right)h\rVert_{X_{\scriptscriptstyle{\mathsf{C}}}}$
@@
$Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})<1\quad\text{and}\quad Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{C}}}^{*})<\frac{(1-Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}}))^{2}}{2Y(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})}$
@@
$r_{\scriptscriptstyle{\mathsf{C}}}=\frac{1-Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})-\sqrt{(1-Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{C}}}^{*}))^{2}-2Y(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{C}}}^{*})}}{Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{C}}}^{*})}.$
@@ 4 Examples of Constructive Proofs of Existence of Gap Solitons
The Gross-Pitaevskii equation (2) with parameters $a=1.1025$, $b=0.55125$, and c=0.826875c=-0.826875 has a soliton solution $u:\mathbb{R}\to\mathbb{R}$, satisfying
@@
$Y(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})=2.6879100002352747\cdot 10^{-13},\quad Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})=0.3465291783592818,\quad Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})=14.980732463866438,$
@@
The second stage of our proof corresponds to the construction of the parameterization of the stable manifold. In this case, for $r_{\scriptscriptstyle{\mathsf{TF}}}^{*}=10^{-3}$ , the bounds $Y(\bar{w},r_{\scriptscriptstyle{\mathsf{F}}})=6.327932449800631\cdot 10^{-9}$, $Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}})=0.9583731072113382$, and $Z_{2}(\bar{w},r_{\scriptscriptstyle{\mathsf{TF}}}^{*})=104.77593347038471$ defined in (50), (52), and (55) satisfy
$Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}})<1\quad\text{and}\quad Z_{2}(\bar{w},r_{\scriptscriptstyle{\mathsf{TF}}}^{*})<\frac{(1-Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}}))^{2}}{2Y(\bar{w},r_{\scriptscriptstyle{\mathsf{F}}})}.$
@@
The third and final stage is to validate a numerical approximation of the corresponding boundary-value problem. For this case, we set $\theta=1$ and $\mathsf{L}=1+2\pi$ in (59). The value for the Taylor variable is $\bar{\sigma}=0.927447198734628.$ As in the previous stages, we fix a norm weight $\omega=1.05$ in (17) and compute the following bounds:
$Y(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})=7.814019760054922\cdot 10^{-7},\quad Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}})=0.9076283031949424\quad$
$\text{and}\quad Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}}^{*})=372.96640912543626,$
@@
$Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})<1\quad\text{and}\quad Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{C}}}^{*})<\frac{\bigl(1-Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})\bigr)^{2}}{2Y(\bar{x}_{\scriptscriptstyle{\mathsf{C}}},r_{\scriptscriptstyle{\mathsf{TF}}})}$
@@
The Gross-Pitaevskii equation (2) with parameters $a=1$, b=1b=-1 and $c=1$, has a soliton solution.